### v1

# LLMs Do Not Have Human-Like Working Memory

Abstract
Human working memory is an active cognitive system that enables not only the temporary storage of information but also its processing and utilization. Without working memory, individuals may produce unreal conversations, exhibit self-contradiction, and struggle with tasks requiring mental reasoning. In this paper, we demonstrate that Large Language Models (LLMs) lack this human-like cognitive ability, posing a significant challenge to achieving artificial general intelligence. We validate this claim through three experiments: (1) Number Guessing Game, (2) Yes or No Game, and (3) Math Magic. Experimental results on several model families indicate that current LLMs fail to exhibit human-like cognitive behaviors in these scenarios. By highlighting this limitation, we aim to encourage further research in developing LLMs with improved working memory capabilities.
Large Language Model,Working Memory
1 Introduction
Imagine the following scenario: You select a number between one and ten. When ready, you are asked, “Is the number greater than five?” Upon answering, others can infer that the number has likely [^1] A possibility remains that your response was made by chance. entered your conscious awareness, as its clear perception is necessary for making the comparison and providing a response.
Figure 1: When ChatGPT says it already has a number in mind, and it is not 4, how can we know whether ChatGPT is lying?
Such conscious awareness is commonly referred to as Working Memory (Atkinson & Shiffrin, 1968). In contrast to long-term memory, working memory is the system required to maintain and manipulate information during complex tasks such as reasoning, comprehension, and learning (Baddeley, 2010). Deficits in working memory can impair information processing and hinder effective communication (Gruszka & Nkecka, 2017; Cowan, 2014).
In this paper, we demonstrate that, despite exhibiting increasingly human-like abilities (Huang et al., 2024b, a), Large Language Models (LLMs) lack a fundamental aspect of human cognition: working memory. As a result, LLMs generate unrealistic responses, display self-contradictions, and fail in tasks requiring mental manipulation.
Previous studies have used the N-back task (Kirchner, 1958) to evaluate the working memory of LLMs (Gong et al., 2024; Zhang et al., 2024). However, a fundamental limitation of these tests is that the critical information needed to answer correctly is in the input context accessible to LLMs. This differs markedly from human testing, where participants cannot view prior steps. LLMs, in contrast, can simply attend to earlier input tokens retained in their context window. To more accurately assess working memory, it is essential to design experiments in which the key information is not explicitly present in the context.
(a) Probabilities for OpenAI GPT models.
(b) Probabilities for OpenAI o-series models.
(c) Probabilities for Meta LLaMA models.
(d) Probabilities for Qwen and DeepSeek models.
Figure 2: Probabilities of model answering “Yes” for each number from one to ten.
The core challenge is to prove that there is actually nothing in LLMs’ mind without knowing what exactly is in their mind. To investigate this limitation, we design three experiments—(1) Number Guessing Game, (2) Yes-No Game, and (3) Math Magic—that test whether LLMs can maintain information internally without explicitly externalizing it in the context. Experimental results reveal that LLMs consistently fail in this capacity, regardless of model family (GPT (Hurst et al., 2024), Qwen (Yang et al., 2024), DeepSeek (Liu et al., 2024; Guo et al., 2025), LLaMA (Grattafiori et al., 2024)) or reasoning approach (Chain-of-Thought (CoT) (Wei et al., 2022) or o1-like reasoning (Jaech et al., 2024)).
2 Number Guessing Game
Consider a number-guessing game in which a human participant privately selects a number between one and ten, and we ask whether the chosen number is one. By repeating this process hundreds of times, we can estimate the probability of the participant selecting the number one. Repeating the procedure for each possible number yields an empirical distribution over the participant’s choices. It is worth noting that, if the participant actually selects a number and responds honestly, the estimated probabilities across all numbers should sum to one.
Setup.
Leveraging this property, we design an experiment to probe LLMs. In each trial, the model is given a fixed prompt: “User: Think of an integer between 1 and 10, but don’t say it to me. Assistant: Got it! I’ve thought of an integer between 1 and 10. What’s next?” The model is then prompted 200 times for each integer from 1 to 10 with queries such as “Is the number you’re thinking of 7? Answer Yes or No.” We record the frequency of “Yes” responses for each number and treat these frequencies as probabilities. If the sum of these probabilities deviates significantly from one, it suggests that LLMs either do not think of a specific number or lie to users. All models are configured with a temperature of 1.0 and a top‑p value of 1.0.
Findings.
Fig. 2 presents the probabilities of “Yes” responses in each model for numbers from one to ten. Two key observations emerge: (1) Most LLMs never produce a “Yes” response; “No” dominates across models. This suggests that LLMs generate unrealistic distributions, implying an inability to produce realistic responses. Given that LLMs generally follow instructions and do not deliberately deceive, we attribute this behavior to their failure to internally “think of” a number. (2) When LLMs do respond affirmatively, they exhibit a marked preference for the number seven. This tendency mirrors patterns observed in humans (Miller, 1956; Kubovy & Psotka, 1976).
Table 1: The sum of probabilities of each model responding “Yes” for all numbers from one to ten. Color intensity reflects proximity to one: red indicates values closer to zero, while blue signifies values greater than one.
Model Sum
GPT-4o-Mini-2024-07-18 0
GPT-4o-2024-05-13 0
GPT-4o-2024-08-06 1.085
GPT-4o-2024-11-20 0
GPT-4.1-2025-04-14 0
o1-Mini-2024-09-12 0.005
o3-Mini-2025-01-31 0.205
o4-Mini-2025-04-16 0.030
LLaMA-3.3-70B-Instruct-Turbo 0.045
LLaMA-3.1-8B-Instruct-Turbo 0.980
LLaMA-3.1-70B-Instruct-Turbo 0.465
LLaMA-3.1-405B-Instruct-Turbo 1.195
Qwen2.5-7B-Instruct-Turbo 0
Qwen2.5-72B-Instruct-Turbo 0
QwQ-32B 0.005
DeepSeek-V3 0
DeepSeek-R1 0.640
Figure 3: Probabilities of GPT-4o-2024-08-06 answering “Yes” for each number in different ranges.
Table 2: The sum of probabilities of GPT-4o-2024-08-06 responding “Yes” for all numbers in different ranges. Color intensity reflects proximity to one: red indicates values closer to zero, while blue signifies values greater than one.
Number Sum
2 0
3 2.180
5 1.600
7 0.920
9 2.025
10 1.085
20 0.080
30 0
40 0.005
We quantify LLM performance on this task using the sum of probabilities. A value closer to one indicates better model performs. Table 1 reports these sums for each model. Several observations emerge: (1) Newer models do not necessarily outperform older ones. Within the GPT family, the 0806 version of GPT-4o (the current target of the “gpt-4o” API) achieves the best performance, surpassing both the 1120 version and GPT-4.1. Similarly, LLaMA-3.3 underperforms relative to LLaMA-3.1. (2) Using CoT reasoning does not improve performance. Models employing such strategies—o1, o3, o4, QwQ, and DeepSeek-R1—fail to produce probability sums closer to one. (3) Overall, LLaMA-3.1 performs best, with the 8B variant outperforming both the 405B and 70B versions.
Other Number Ranges.
We extend our experiments to include a broader range of numbers. Given that GPT-4o-2024-08-06 performs best among the OpenAI models, we focus on its behavior across different numerical ranges. Table 2 reports the summed probabilities for each range, while Fig. 3 illustrates the probability of individual numbers. Our findings reveal two key patterns: (1) For smaller ranges such as 3, 5, and 9, the model exhibits a strong bias toward answering “Yes,” with the probability sum significantly exceeding one. In contrast, for larger ranges like 20, 30, and 40, “Yes” responses are rare. (2) When the model does produce a “Yes” response, it frequently corresponds to numbers ending in seven (e.g., 7, 17, 37), as shown in Fig. 3.
In conclusion, LLMs fail to generate correct distributions, suggesting they struggle to produce realistic responses. This limitation likely arises from their inability to internally think of numerical values without explicit contextual cues.
3 Yes-No Game
“Yes-No” is a social deduction game commonly used to train human reasoning, classification, and questioning skills. In this game, one player privately selects an object, while the opponent asks yes–no questions (e.g., “Is the object heavier than an elephant?”) to progressively narrow the possibilities and ultimately guess the object. Consider the decision-making process of the player answering questions. Upon receiving a question, the player simply compares the imagined object with the queried attribute and responds accordingly. Note that, humans do not typically recall all previous questions and answers to ensure consistency; instead, they answer each new question based on immediate judgment, without explicitly checking self-contradiction with prior responses.
We hypothesize that if LLMs lack the working memory to temporarily retain an imagined object, they can only respond to questions by checking consistency with their prior answers. As the number of questions increases, maintaining consistency becomes increasingly difficult. Consequently, performance on this task is closely tied to the model’s ability for long-context reasoning. In this section, we present an experiment in which the LLM is first instructed to imagine an object, followed by a sequence of comparative questions involving that object and other reference objects. The goal is to assess whether the model produces self-contradictions when answering sequential questions about the imagined object. For instance, the model might initially answer “Yes” to “Is the object heavier than an elephant?” but later also respond “Yes” to “Is the object lighter than a cat?”, thereby contradicting itself.
Setup.
We predefined five sets of objects that are commonly regarded as comparable with respect to five properties: volume, length, weight, density, and hardness. In total, 60 distinct objects are included, as listed in Table 3, ordered by the corresponding property. For each question, one property is randomly selected, followed by an object from the corresponding object list. The LLM is then prompted to assess whether the object it has in mind is comparative than the given object, where the comparative form can vary in direction (e.g., bigger or smaller for volume). In each trial, the model is continuously presented with up to 250 such questions. We record the number of questions completed before the model exhibits a self-contradiction. If no contradiction is observed across all 250 questions, the trial is considered a Pass. Each model is tested with 200 trials.
Table 3: Objects ordered by the five properties (smallest to largest).
Volume Length Weight Density
Coffee bean Rice Coin Air
Dice Paperclip Spoon Wood
Golf ball Credit card Watch Ice
Soda can Pencil Smartphone Water
Soccer ball Laptop Bottle of water Plastic
Microwave oven Baseball bat Dictionary Glass
Washing machine Guitar Cat Iron
Bathtub Door Bicycle Copper
Car Apple tree Television Silver
School bus Coconut tree Refrigerator Gold
Shipping container Tennis court Tiger Hardness
Olympic swimming pool Swimming pool Cow Marshmallow
Boeing 747 Football field Rhino Rubber eraser
Titanic Skyscraper Elephant Brick
Great Pyramid of Giza Mount Everest Train Hammer
Diamond ring
Table 4: Count of failures of GPT-4o-Mini-2024-07-18 and GPT-4o-2024-08-06 on different properties.
Model Failure V W L D H
GPT-4o-Mini 200 12 46 49 52 41
GPT-4o 173 21 42 57 27 26
Findings.
Table 4 presents the number of failed trials for the two models. The smaller model (GPT-4o-Mini) consistently fails, indicating that it always produces self-contradiction. As model capacity increases, GPT-4o successfully passes 27 out of 200 trials. This result supports our hypothesis that model performance on this task depends on the their long-context processing ability—an outcome that would be unlikely if the models possessed the working memory for maintaining imagined objects.
Figure 4 presents histograms of the number of questions each model completes before exhibiting self-contradiction. The distribution for GPT-4o-Mini peaks in the 20–30 range, whereas GPT-4o peaks in the 30–40 range. Moreover, GPT-4o demonstrates a higher frequency of completions in the 80–130 range compared to GPT-4o-Mini. Notably, the types of properties that lead to self-contradictions differ between the two models: GPT-4o-Mini more frequently fails on density and hardness, while GPT-4o shows greater robustness on these attributes.
In conclusion, LLMs exhibit self-contradictions (e.g., claiming an object is both larger than a car and smaller than a soccer ball) as the number of queries increases. This behavior suggests their reliance on long-context reasoning rather than a proper working memory.
(a) GPT-4o-Mini-2024-07-18.
(b) GPT-4o-2024-08-06.
Figure 4: The histogram of the number of questions where the two models show self-contradiction.
4 Math Magic
Let’s play a game before presenting the final part of our experiments. Think of a three-digit number in which the hundreds and units digits differ (e.g., abc). Reverse the digits to form a new number (e.g., abc $\rightarrow$ cba), and subtract the smaller of the two numbers from the larger. Then, reverse the result—if it is a two-digit number, prepend a zero (e.g., 67 $\rightarrow$ 067 $\rightarrow$ 760). Add this reversed number to the previous result What if we can tell that your result is 1089? It is math who tells us the secret. Such magic, often referred to as “math magic” or “math mentalism,” rely on carefully constructed transformations that yield predictable outcomes, allowing magicians to perform mind-reading.
To perform this task, humans must internally generate or retain numerical information (e.g., remember a poker card) and manipulate it mentally without external cues. This process relies on working memory, a cognitive function that LLMs may lack. Without this capacity, LLMs are prone to errors in mathematical tricks, failing to produce the intended outcomes. Accordingly, we can design complex math magics to evaluate whether LLMs possess working memory–like capabilities.
Table 5: Operations in the math magic in our experiment. the random number $a$ ranges from 1 to 7, while the random numbers $b$ and $c$ range from 1 to 3.
Role Content
User Think of 4 integers between 1 and {NUMBER} in order, but don’t tell me.
Assistant Okay! I’ve got 4 numbers. What’s next?
User In order, append the same 4 numbers after the original ones.
Assistant Understood! Now I have 8 numbers. What’s next?
User Take the first {random_number_a} numbers and move them to the end.
Assistant Got it! Now I have moved the numbers. What’s next?
User Take the first 3 numbers and insert them anywhere in the middle.
Assistant Okay! The first 3 numbers are placed somewhere in the middle. What’s next?
User Set the first number aside. We don’t need it for now.
Assistant Understood! Now I have 7 numbers. What’s next?
User Take the first {random_number_b} numbers and insert them anywhere in the middle.
Assistant Got it. The first {random_number_b} numbers are placed somewhere in the middle. What’s next?
User Remove the first {random_number_c} numbers. We will never need it anymore.
Assistant Okay! Now I have {8 - random_number_c} numbers. What’s next?
User Take the first number and move it to the end. Repeat this seven times.
Assistant Understood! Now my sequence has rearranged. What’s next?
User Take the first number and move it to the end, then remove the second number. Repeat this {7 - random_number_c} times.
Assistant Got it! Now I have only 1 number. What’s next?
User Tell me what the last remaining number is. Do you remember the number you set aside at the beginning? Tell me what that number was.
Setup.
Since the 1089 game is well known and likely included in the training data of LLMs, our preliminary experiments show that LLMs can recognize the game and thus invalidate the evaluation. Therefore, to more effectively assess LLMs’ multi-step mental manipulation, we select a more complicated math magicbased on the Josephus Problem (Schumer, 2002). In this task, participants are asked to imagine four numbers and perform a sequence of operations, including duplication, rotation, and removal. The procedure is shown in Table 5. Ultimately, only two numbers remain, and mathematical constraints guarantee they are identical. In our experiment, LLMs are prompted to secretly think of four numbers and mentally execute the sequence of randomized operations. We measure the proportion of trials (out of 150) in which the LLM completes the task correctly.
Table 6: Model performance on the Math Magic. GPT-4.1-2025-04-14 fails to complete the task, incorrectly assuming that the necessary numerical inputs are missing.
Model Count Acc (%)
GPT-4o-Mini-2024-07-18 0/150 0.0
GPT-4o-2024-05-13 6/150 4.0
GPT-4o-2024-08-06 7/150 4.7
GPT-4o-2024-11-20 1/150 0.7
GPT-4.1-2025-04-14 - -
LLaMA-3.3-70B-Instruct-Turbo 21/150 14.0
LLaMA-3.1-8B-Instruct-Turbo 33/150 22.0
LLaMA-3.1-70B-Instruct-Turbo 12/150 8.0
LLaMA-3.1-405B-Instruct-Turbo 28/150 18.7
Qwen2.5-7B-Instruct-Turbo 0/150 0.0
Qwen2.5-72B-Instruct-Turbo 8/150 5.3
DeepSeek-V3 0/150 0.0
Table 7: Performance of LLMs with CoT or LRMs on the Math Magic. Similar to GPT-4.1-2025-04-14, o4-Mini-2025-04-16 also fails to complete the task.
Model w/ CoT or LRM Count Acc (%)
GPT-4o-Mini-2024-07-18 3/150 2.0
GPT-4o-2024-08-06 32/150 21.3
LLaMA-3.3-70B-Instruct-Turbo 33/150 22.0
Qwen2.5-72B-Instruct-Turbo 23/150 15.3
DeepSeek-V3 18/150 12.0
o1-Mini-2024-09-12 25/150 16.7
o3-Mini-2025-01-31 3/150 2.0
o4-Mini-2025-04-16 - -
QwQ-32B 45/150 30.0
DeepSeek-R1 59/150 39.3
Findings.
Table 6 reports the accuracy of prompting models to output the two numbers directly. Most LLMs perform poorly on this task, with notably higher accuracy observed in the LLaMA model family. This finding aligns with results from the number guessing game shown in Table 1, where LLaMA models generate more realistic distributions than other models. Together, these consistent results suggest that current LLMs generally lack the working memory required for such tasks.
We also investigate whether CoT prompting can enhance performance on this task. Table 7 presents the results of prompting models to reason step by step, as well as the performance of o1-like long reasoning models (LRMs). The best-performing model is DeepSeek-R1, achieving an accuracy of 39.3%. Most models attain only 10–20% accuracy—marginally above the random guessing baseline of 10%—highlighting substantial room for improvement. Notably, as observed in the number guessing experiment, models exhibit a preference for the number seven, which appears to influence performance here. Of the 59 correct predictions by DeepSeek-R1, 48 (81.4%) involve selecting the number 7; for o1-Mini, the number is 22 out of 25 (88%). These results suggest that models may not fully comprehend the task and instead rely on biased number preferences.
In conclusion, current LLMs struggle with a class of tasks that remain challenging even with CoT prompting. These tasks demand internal object imagination and mental manipulation, relying on an effective working memory.
5 Related Work
Working memory has been widely discussed in the context of LLMs. However, most studies use the term to describe LLMs’ capacity to process information within a fixed context window (Li et al., 2023; Guo et al., 2023). Gong et al. (2024); Zhang et al. (2024) evaluate LLM working memory using N-back tasks. As noted in §1, however, such human-designed tests may not be valid for LLMs.
A separate line of work defines LLM working memory as an explicitly designed external module for LLM agents (Hu et al., 2024; Zeng et al., 2024). Wang et al. (2024) incorporate working memory to enhance reasoning, while Kang et al. (2024) use it to improve training efficiency. These approaches, however, do not explore whether LLMs possess a form of human-like working memory—such as the ability to mentally simulate and manipulate objects.
6 Conclusion
In this study, we present three carefully designed experiments to investigate whether LLMs possess human-like working memory for processing information and generating responses. Across all experiments, the results reveal a consistent pattern: LLMs do not exhibit behavior indicative of a functional working memory. They fail to internally represent and manipulate transient information across multiple reasoning steps, relying instead on the immediate prompt context. Even advanced prompting strategies, such as CoT prompting, yield only marginal improvements on tasks requiring internal state management. This limitation leads to unrealistic responses, self-contradictions, and an inability to perform mental manipulations. We hope these findings encourage future research on integrating working memory mechanisms into LLMs to enhance their reasoning capabilities and human-likeness.

### v2

@@
-LLMs Do Not Have Human-Like Working Memory
+Language Models Do Not Have Human-Like Working Memory
 Abstract
+While Large Language Models (LLMs) exhibit remarkable reasoning abilities, we demonstrate that they fundamentally lack a core aspect of human cognition: working memory.
 Human working memory is an active cognitive system that enables not only the temporary storage of information but also its processing and utilization.
 Without working memory, individuals may produce unreal conversations, exhibit self-contradiction, and struggle with tasks requiring mental reasoning.
-In this paper, we demonstrate that Large Language Models (LLMs) lack this human-like cognitive ability, posing a significant challenge to achieving artificial general intelligence.
-We validate this claim through three experiments: (1) Number Guessing Game, (2) Yes or No Game, and (3) Math Magic.
-Experimental results on several model families indicate that current LLMs fail to exhibit human-like cognitive behaviors in these scenarios.
-By highlighting this limitation, we aim to encourage further research in developing LLMs with improved working memory capabilities.
-Large Language Model,Working Memory
+Existing evaluations using N-back or context-dependent tasks fail as they allow LLMs to exploit accessible context rather than retain latent information.
+We introduce three novel tasks—(1) Number Guessing, (2) Yes-No Deduction, and Math Magic—that isolate internal representation from external context.
+Across seventeen frontier models spanning four major model families, we consistently observe irrational or contradictory behaviors, highlighting LLMs’ inability to retain and manipulate latent information.
+Our work establishes a new benchmark for evaluating working memory in LLMs and identifies this deficit as a critical obstacle to artificial general intelligence.
+Code and prompts for the experiments are available at https://github.com/penguinnnnn/LLM-Working-Memory.
+Figure 1: When LLMs say they already have a number in mind, and it is not 4, how can we know whether LLMs are lying, or even thinking of nothing?
 1 Introduction
+Large Language Models (LLMs) have demonstrated strong performance across diverse domains, including legal advice Guha et al.
+(2023), education Wen et al.
+(2024), translation Jiao et al.
+(2023), and healthcare Yang et al.
+(2024b).
+Beyond task performance, recent studies have examined their intrinsic behaviors, often drawing parallels with human cognition, such as personality Huang et al.
+(2024b), emotion Huang et al.
+(2024a), empathy Sorin et al.
+(2024), theory-of-mind Liu et al.
+(2024b), and values Wang et al.
+(2024b).
+One such behavior is memory, which has attracted increasing attention from both industry and research communities.
+OpenAI was the first to introduce a memory module in ChatGPT (February 2024).
+[^1] https://openai.com/index/memory-and-new-controls-for-chatgpt/ By mid-2025, xAI, Anthropic, and Google integrated memory into Grok, [^2] https://x.com/grok/status/1912670182012801156 Claude, [^3] https://www.anthropic.com/news/claude-4 and Gemini, [^4] https://blog.google/products/gemini/temporary-chats-privacy-controls/ respectively.
+Sam Altman even emphasized that “people want memory,” identifying it as a central feature of the forthcoming GPT-6.555https://futurism.com/disastrous-gpt-5-sam-altman-hyping-up-gpt-6 On the research side, memory modules are also actively studied Zhang et al.
+(2025); Hatalis et al.
+(2023); Xu et al.
+(2025).
+What is memory?
+Atkinson & Shiffrin (1968) categorized memory by retention timescale: sensory (1ms–2s), short-term (seconds), and long-term (hours to lifetime).
+Long-term memory enables information storage over extended periods, whereas short-term, or Working Memory, maintains and manipulates information during complex tasks such as reasoning, comprehension, and learning Baddeley & Hitch (1974).
+These distinctions have also been adopted in machine learning, corresponding to representational learning for raw inputs (sensory memory), in-context computation at test time (working memory), and access to external databases (long-term memory) Weng (2023).
+As a central aspect of human cognition, memory has been widely benchmarked Du et al.
+(2025).
+However, working memory remains relatively underexplored.
+Existing studies often adopt the N-back task Kirchner (1958) to assess LLMs’ working memory Gong et al.
+(2024); Zhang et al.
+(2024).
+Yet, a fundamental limitation arises: the critical information for correct responses remains accessible in the model’s input context.
+Unlike humans, who cannot revisit prior steps, LLMs can simply attend to earlier tokens within their context window.
+Fig.
+2 illustrates this discrepancy.
+To more faithfully evaluate working memory, it is necessary to design experiments where the key information is not explicitly present in the context.
 Imagine the following scenario: You select a number between one and ten.
-When ready, you are asked, “Is the number greater than five?” Upon answering, others can infer that the number has likely [^1] A possibility remains that your response was made by chance.
-entered your conscious awareness, as its clear perception is necessary for making the comparison and providing a response.
-Figure 1: When ChatGPT says it already has a number in mind, and it is not 4, how can we know whether ChatGPT is lying?
-Such conscious awareness is commonly referred to as Working Memory (Atkinson & Shiffrin, 1968).
-In contrast to long-term memory, working memory is the system required to maintain and manipulate information during complex tasks such as reasoning, comprehension, and learning (Baddeley, 2010).
-Deficits in working memory can impair information processing and hinder effective communication (Gruszka & Nkecka, 2017; Cowan, 2014).
-In this paper, we demonstrate that, despite exhibiting increasingly human-like abilities (Huang et al., 2024b, a), Large Language Models (LLMs) lack a fundamental aspect of human cognition: working memory.
-As a result, LLMs generate unrealistic responses, display self-contradictions, and fail in tasks requiring mental manipulation.
-Previous studies have used the N-back task (Kirchner, 1958) to evaluate the working memory of LLMs (Gong et al., 2024; Zhang et al., 2024).
-However, a fundamental limitation of these tests is that the critical information needed to answer correctly is in the input context accessible to LLMs.
-This differs markedly from human testing, where participants cannot view prior steps.
-LLMs, in contrast, can simply attend to earlier input tokens retained in their context window.
-To more accurately assess working memory, it is essential to design experiments in which the key information is not explicitly present in the context.
+When ready, you are asked, “Is the number greater than five?” If you answer, observers can reasonably infer that the number has entered your conscious awareness (i.e., your working memory), since clear perception is necessary to perform the comparison and provide a response.
+[^6] A possibility remains that the response was given by chance.
+We ask: Do LLMs possess human-like working memory beyond evaluations that merely test their ability to attend to visible information?
+AI models are often viewed as reasoning in two modes: (1) the token space over sequences Wei et al.
+(2022); Yao et al.
+(2023), and (2) the latent space over activations Hao et al.
+(2024); Geiping et al.
+(2025).
+Our focus is on the second mode, where evaluating working memory provides insight into whether models can hold and manipulate latent concepts without explicit externalization.
+Success in this capacity could enhance reasoning without reliance on Chain-of-Thought (CoT), as it directly tests the model’s ability to maintain objects and concepts internally.
+Conversely, deficits in working memory impair information processing in humans Gruszka & Nkecka (2017); Cowan (2014), and in LLMs manifest as unrealistic outputs, self-contradictions, and failures on tasks requiring mental manipulation.
+The central challenge in designing such evaluations is epistemic: how can we demonstrate the absence of internal memory when we cannot directly observe a model’s mind?
+To address this limitation, we design three experiments—(1) Number Guessing, (2) Yes–No Deduction, and (3) Math Magic—that test whether LLMs can internally maintain information that is not explicitly present in the input context.
+Our experiments span 17 state-of-the-art LLMs, both proprietary and open-source, and reveal a consistent inability to perform such tasks.
+The failure is uniform across model families (GPT Hurst et al.
+(2024), Qwen Yang et al.
+(2024a), DeepSeek Liu et al.
+(2024a); Guo et al.
+(2025), LLaMA Grattafiori et al.
+(2024)) and reasoning approaches (CoT Wei et al.
+(2022) or o1-like reasoning Jaech et al.
+(2024)).
+2 Background
+Figure 2: a.
+An illustration of how “N-Back” tasks are performed.
+b.
+Humans see the stimuli one after one, forcing them to put the information in working memory.
+c.
+Researchers put all stimuli into context, enabling LLMs to easily find the answers.
+2.1 How Working Memory is Evaluated
+Human working memory is typically assessed using behavioral paradigms that require individuals to maintain and update information over short intervals.
+Common examples include the digit span task Miller (1956), where participants recall sequences of numbers of increasing length, and the N-back task Kirchner (1958), which requires identifying whether the current stimulus matches one presented $N$ steps earlier.
+These tasks are widely used because they probe the ability to maintain and manipulate information that is no longer externally visible, thereby capturing the essence of working memory function.
+In the context of LLMs, working memory has been widely discussed.
+However, most studies use the term to describe an LLM’s capacity to process information within a fixed context window Li et al.
+(2023); Guo et al.
+(2023).
+Gong et al.
+(2024); Zhang et al.
+(2024) evaluate LLMs using N-back tasks.
+As noted in §1, such human-designed tests are not directly valid for LLMs, since models can simply attend to retained tokens in their context window rather than actively maintaining information.
+2.2 Existing Engineering Solutions
+A separate line of work introduces explicit engineering solutions by equipping LLMs with external memory modules Hu et al.
+(2024); Zeng et al.
+(2024).
+For example, Wang et al.
+(2024a) incorporate symbolic working memory to enhance reasoning, while Kang et al.
+(2024) use it to improve training efficiency.
+Other approaches implement scratch spaces Lanchantin et al.
+(2023) or internal CoT mechanisms Jaech et al.
+(2024), which may partially mitigate working memory limitations identified in this paper.
+However, these methods do not address whether LLMs possess an intrinsic working memory capacity, analogous to humans’ ability to mentally simulate and manipulate objects.
+The distinction can be illustrated with an analogy: modern calculators make arithmetic trivial, yet schools continue to assess addition, subtraction, multiplication, and division.
+The purpose is not solely to solve problems efficiently, but to reveal individuals’ underlying cognitive abilities.
+Similarly, while engineering techniques can extend an LLM’s effective memory, we are ultimately interested in whether the model itself has the basic cognitive capability of working memory.
+Without it, a model may function adequately through external tools, but its intrinsic reasoning ability remains fundamentally limited.
+3 Number Guessing
 (a) Probabilities for OpenAI GPT models.
@@
 Figure 3: Probabilities of model answering “Yes” for each number from one to ten.
-The core challenge is to prove that there is actually nothing in LLMs’ mind without knowing what exactly is in their mind.
-To investigate this limitation, we design three experiments—(1) Number Guessing Game, (2) Yes-No Game, and (3) Math Magic—that test whether LLMs can maintain information internally without explicitly externalizing it in the context.
-Experimental results reveal that LLMs consistently fail in this capacity, regardless of model family (GPT (Hurst et al., 2024), Qwen (Yang et al., 2024), DeepSeek (Liu et al., 2024; Guo et al., 2025), LLaMA (Grattafiori et al., 2024)) or reasoning approach (Chain-of-Thought (CoT) (Wei et al., 2022) or o1-like reasoning (Jaech et al., 2024)).
-2 Number Guessing Game
 Consider a number-guessing game in which a human participant privately selects a number between one and ten, and we ask whether the chosen number is one.
@@
 It is worth noting that, if the participant actually selects a number and responds honestly, the estimated probabilities across all numbers should sum to one.
+Figure 4: Probabilities of GPT-4o-2024-08-06 answering “Yes” for each number in different ranges.
 Setup.
@@
 All models are configured with a temperature of 1.0 and a top‑p value of 1.0.
-Findings.
-Fig.
-2 presents the probabilities of “Yes” responses in each model for numbers from one to ten.
-Two key observations emerge: (1) Most LLMs never produce a “Yes” response; “No” dominates across models.
-This suggests that LLMs generate unrealistic distributions, implying an inability to produce realistic responses.
-Given that LLMs generally follow instructions and do not deliberately deceive, we attribute this behavior to their failure to internally “think of” a number.
-(2) When LLMs do respond affirmatively, they exhibit a marked preference for the number seven.
-This tendency mirrors patterns observed in humans (Miller, 1956; Kubovy & Psotka, 1976).
 Table 1: The sum of probabilities of each model responding “Yes” for all numbers from one to ten.
@@
 DeepSeek-R1 0.640
-Figure 3: Probabilities of GPT-4o-2024-08-06 answering “Yes” for each number in different ranges.
+Findings.
+Fig.
+3 presents the probabilities of “Yes” responses in each model for numbers from one to ten.
+Two key observations emerge: (1) Most LLMs never produce a “Yes” response; “No” dominates across models.
+This suggests that LLMs generate unrealistic distributions, implying an inability to produce realistic responses.
+Given that LLMs generally follow instructions and do not deliberately deceive, we attribute this behavior to their failure to internally “think of” a number.
+(2) When LLMs do respond affirmatively, they exhibit a marked preference for the number seven.
+This tendency mirrors patterns observed in humans Miller (1956); Kubovy & Psotka (1976).
+We quantify LLM performance on this task using the sum of probabilities.
+A value closer to one indicates better model performs.
+Table 1 reports these sums for each model.
+Several observations emerge: (1) Newer models do not necessarily outperform older ones.
+Within the GPT family, the 0806 version of GPT-4o (the current target of the “gpt-4o” API) achieves the best performance, surpassing both the 1120 version and GPT-4.1.
+Similarly, LLaMA-3.3 underperforms relative to LLaMA-3.1.
+(2) Using CoT reasoning does not improve performance.
+Models employing such strategies—o1, o3, o4, QwQ, and DeepSeek-R1—fail to produce probability sums closer to one.
+(3) Overall, LLaMA-3.1 performs best, with the 8B variant outperforming both the 405B and 70B versions.
 Table 2: The sum of probabilities of GPT-4o-2024-08-06 responding “Yes” for all numbers in different ranges.
-Color intensity reflects proximity to one: red indicates values closer to zero, while blue signifies values greater than one.
 Number Sum
@@
 40 0.005
-We quantify LLM performance on this task using the sum of probabilities.
-A value closer to one indicates better model performs.
-Table 1 reports these sums for each model.
-Several observations emerge: (1) Newer models do not necessarily outperform older ones.
-Within the GPT family, the 0806 version of GPT-4o (the current target of the “gpt-4o” API) achieves the best performance, surpassing both the 1120 version and GPT-4.1.
-Similarly, LLaMA-3.3 underperforms relative to LLaMA-3.1.
-(2) Using CoT reasoning does not improve performance.
-Models employing such strategies—o1, o3, o4, QwQ, and DeepSeek-R1—fail to produce probability sums closer to one.
-(3) Overall, LLaMA-3.1 performs best, with the 8B variant outperforming both the 405B and 70B versions.
 Other Number Ranges.
@@
 (2) When the model does produce a “Yes” response, it frequently corresponds to numbers ending in seven (e.g., 7, 17, 37), as shown in Fig.
-3.
+4.
 In conclusion, LLMs fail to generate correct distributions, suggesting they struggle to produce realistic responses.
 This limitation likely arises from their inability to internally think of numerical values without explicit contextual cues.
-3 Yes-No Game
+4 Yes-No Deduction
+Table 3: Objects ordered by the five properties (smallest to largest).
+Volume Length Weight Density
+Coffee bean Rice Coin Air
+Dice Paperclip Spoon Wood
+Golf ball Credit card Watch Ice
+Soda can Pencil Smartphone Water
+Soccer ball Laptop Bottle of water Plastic
+Microwave oven Baseball bat Dictionary Glass
+Washing machine Guitar Cat Iron
+Bathtub Door Bicycle Copper
+Car Apple tree Television Silver
+School bus Coconut tree Refrigerator Gold
+Shipping container Tennis court Tiger Hardness
+Olympic swimming pool Swimming pool Cow Marshmallow
+Boeing 747 Football field Rhino Rubber eraser
+Titanic Skyscraper Elephant Brick
+Great Pyramid of Giza Mount Everest Train Hammer
+Diamond ring
 “Yes-No” is a social deduction game commonly used to train human reasoning, classification, and questioning skills.
@@
 Each model is tested with 200 trials.
-Table 3: Objects ordered by the five properties (smallest to largest).
-Volume Length Weight Density
-Coffee bean Rice Coin Air
-Dice Paperclip Spoon Wood
-Golf ball Credit card Watch Ice
-Soda can Pencil Smartphone Water
-Soccer ball Laptop Bottle of water Plastic
-Microwave oven Baseball bat Dictionary Glass
-Washing machine Guitar Cat Iron
-Bathtub Door Bicycle Copper
-Car Apple tree Television Silver
-School bus Coconut tree Refrigerator Gold
-Shipping container Tennis court Tiger Hardness
-Olympic swimming pool Swimming pool Cow Marshmallow
-Boeing 747 Football field Rhino Rubber eraser
-Titanic Skyscraper Elephant Brick
-Great Pyramid of Giza Mount Everest Train Hammer
-Diamond ring
-Table 4: Count of failures of GPT-4o-Mini-2024-07-18 and GPT-4o-2024-08-06 on different properties.
+Table 4: Count of failures of GPT-4o-Mini-2024-07-18 and GPT-4o-2024-08-06.
 Model Failure V W L D H
@@
 Notably, the types of properties that lead to self-contradictions differ between the two models: GPT-4o-Mini more frequently fails on density and hardness, while GPT-4o shows greater robustness on these attributes.
+Table 5: Ablation studies using GPT-4o-2024-08-06.
+Model Failure V W L D H
+Hint 194 37 39 60 37 21
+All 158 18 29 21 55 35
+Hint + All 145 13 32 34 46 20
+Ablation.
+To ensure that the observed failures are not simply due to LLMs’ inability to rank objects by the five properties, we conduct the following ablation studies: (1) Hint: At the beginning of the prompt, we provide GPT-4o with the object rankings defined in Table 3.
+(2) All: For each question, we specify the target object $\mathcal{O}$ by stating explicitly that “the object you are considering is $\mathcal{O}$.” (3) Hint + All: We combine the above two settings.
+Results are shown in Table 5.
+Two key findings emerge: (1) Even when given hints that guarantee knowledge of the correct ranking, GPT-4o still frequently fails, suggesting that the task critically depends on long-context reasoning.
+(2) When the object identity is provided explicitly, error rates decrease, likely because the long-context reasoning task is reduced to a short-context one.
 In conclusion, LLMs exhibit self-contradictions (e.g., claiming an object is both larger than a car and smaller than a soccer ball) as the number of queries increases.
@@
 Assistant Understood! Now I have 8 numbers. What’s next?
-User Take the first {random_number_a} numbers and move them to the end.
+User Move the first {random_number_a} numbers to the end.
 Assistant Got it! Now I have moved the numbers. What’s next?
@@
 User Remove the first {random_number_c} numbers. We will never need it anymore.
-Assistant Okay! Now I have {8 - random_number_c} numbers. What’s next?
-User Take the first number and move it to the end. Repeat this seven times.
+Assistant Okay! Now I have {7 - random_number_c} numbers. What’s next?
+User Move the first number to the end. Repeat this seven times.
 Assistant Understood! Now my sequence has rearranged. What’s next?
-User Take the first number and move it to the end, then remove the second number. Repeat this {7 - random_number_c} times.
+User Remove the second number, and then move the first number to the end. Repeat this {6 - random_number_c} times.
 Assistant Got it! Now I have only 1 number. What’s next?
@@
 Since the 1089 game is well known and likely included in the training data of LLMs, our preliminary experiments show that LLMs can recognize the game and thus invalidate the evaluation.
-Therefore, to more effectively assess LLMs’ multi-step mental manipulation, we select a more complicated math magicbased on the Josephus Problem (Schumer, 2002).
+Therefore, to more effectively assess LLMs’ multi-step mental manipulation, we select a more complicated math magicbased on the Josephus Problem Schumer (2002).
 In this task, participants are asked to imagine four numbers and perform a sequence of operations, including duplication, rotation, and removal.
@@
 We measure the proportion of trials (out of 150) in which the LLM completes the task correctly.
-Table 6: Model performance on the Math Magic. GPT-4.1-2025-04-14 fails to complete the task, incorrectly assuming that the necessary numerical inputs are missing.
+Table 7: LLM performance on the math magic task.
 Model Count Acc (%)
 GPT-4o-Mini-2024-07-18 0/150 0.0
-GPT-4o-2024-05-13 6/150 4.0
-GPT-4o-2024-08-06 7/150 4.7
-GPT-4o-2024-11-20 1/150 0.7
+GPT-4o-2024-05-13 4/150 2.7
+GPT-4o-2024-08-06 3/150 2.0
+GPT-4o-2024-11-20 0/150 0.0
 GPT-4.1-2025-04-14 - -
-LLaMA-3.3-70B-Instruct-Turbo 21/150 14.0
-LLaMA-3.1-8B-Instruct-Turbo 33/150 22.0
-LLaMA-3.1-70B-Instruct-Turbo 12/150 8.0
-LLaMA-3.1-405B-Instruct-Turbo 28/150 18.7
-Qwen2.5-7B-Instruct-Turbo 0/150 0.0
-Qwen2.5-72B-Instruct-Turbo 8/150 5.3
-DeepSeek-V3 0/150 0.0
-Table 7: Performance of LLMs with CoT or LRMs on the Math Magic. Similar to GPT-4.1-2025-04-14, o4-Mini-2025-04-16 also fails to complete the task.
+LLaMA-3.3-70B-Instruct-Turbo 7/150 5.7
+LLaMA-3.1-8B-Instruct-Turbo 20/150 13.3
+LLaMA-3.1-70B-Instruct-Turbo 7/150 5.7
+LLaMA-3.1-405B-Instruct-Turbo 39/150 26.0
+Qwen2.5-7B-Instruct-Turbo 8/150 5.3
+Qwen2.5-72B-Instruct-Turbo 2/150 1.3
+DeepSeek-V3 4/150 2.7
+(a) LLMs without CoT. GPT-4.1-2025-04-14 fails to complete most of the cases, incorrectly assuming that the necessary numerical inputs are missing.
 Model w/ CoT or LRM Count Acc (%)
-GPT-4o-Mini-2024-07-18 3/150 2.0
-GPT-4o-2024-08-06 32/150 21.3
-LLaMA-3.3-70B-Instruct-Turbo 33/150 22.0
-Qwen2.5-72B-Instruct-Turbo 23/150 15.3
-DeepSeek-V3 18/150 12.0
-o1-Mini-2024-09-12 25/150 16.7
-o3-Mini-2025-01-31 3/150 2.0
-o4-Mini-2025-04-16 - -
-QwQ-32B 45/150 30.0
-DeepSeek-R1 59/150 39.3
+GPT-4o-Mini-2024-07-18 5/150 3.3
+GPT-4o-2024-05-13 26/150 17.3
+GPT-4o-2024-08-06 31/150 20.7
+GPT-4o-2024-11-20 - -
+LLaMA-3.3-70B-Instruct-Turbo 25/150 16.7
+Qwen2.5-7B-Instruct-Turbo 49/150 32.7
+Qwen2.5-72B-Instruct-Turbo 37/150 24.7
+DeepSeek-V3 48/150 32.0
+o1-Mini-2024-09-12 75/150 50.0
+o3-Mini-2025-01-31 145/150 96.7
+o4-Mini-2025-04-16 54/150 36.0
+QwQ-32B 135/150 90.0
+DeepSeek-R1 150/150 100
+(b) LLMs with CoT and Large Reasoning Models. GPT-4o-2024-11-20 consistently fails this task.
 Findings.
-Table 6 reports the accuracy of prompting models to output the two numbers directly.
+Table 7a reports the accuracy of prompting models to output the two numbers directly.
 Most LLMs perform poorly on this task, with notably higher accuracy observed in the LLaMA model family.
@@
 Together, these consistent results suggest that current LLMs generally lack the working memory required for such tasks.
-We also investigate whether CoT prompting can enhance performance on this task.
-Table 7 presents the results of prompting models to reason step by step, as well as the performance of o1-like long reasoning models (LRMs).
-The best-performing model is DeepSeek-R1, achieving an accuracy of 39.3%.
-Most models attain only 10–20% accuracy—marginally above the random guessing baseline of 10%—highlighting substantial room for improvement.
-Notably, as observed in the number guessing experiment, models exhibit a preference for the number seven, which appears to influence performance here.
-Of the 59 correct predictions by DeepSeek-R1, 48 (81.4%) involve selecting the number 7; for o1-Mini, the number is 22 out of 25 (88%).
-These results suggest that models may not fully comprehend the task and instead rely on biased number preferences.
-In conclusion, current LLMs struggle with a class of tasks that remain challenging even with CoT prompting.
-These tasks demand internal object imagination and mental manipulation, relying on an effective working memory.
-5 Related Work
-Working memory has been widely discussed in the context of LLMs.
-However, most studies use the term to describe LLMs’ capacity to process information within a fixed context window (Li et al., 2023; Guo et al., 2023).
-Gong et al.
-(2024); Zhang et al.
-(2024) evaluate LLM working memory using N-back tasks.
-As noted in §1, however, such human-designed tests may not be valid for LLMs.
-A separate line of work defines LLM working memory as an explicitly designed external module for LLM agents (Hu et al., 2024; Zeng et al., 2024).
-Wang et al.
-(2024) incorporate working memory to enhance reasoning, while Kang et al.
-(2024) use it to improve training efficiency.
-These approaches, however, do not explore whether LLMs possess a form of human-like working memory—such as the ability to mentally simulate and manipulate objects.
+We further examine whether CoT prompting improves performance on this task.
+Table 7b presents the results of prompting models to reason step by step, as well as the performance of o1-like long reasoning models (LRMs).
+Base models prompted to reason step-by-step achieve 10–30% accuracy—substantially higher than without CoT.
+DeepSeek-R1 attains 100% accuracy, and other LRMs also perform well, indicating that extended token-level reasoning substantially improves accuracy.
+However, consistent with our number-guessing experiment, models exhibit a strong preference for the number seven, which appears to influence performance—especially within the o-series.
+Among o1-Mini’s 75 correct predictions, 50 (66.7%) select 7; for o3-Mini and o4-Mini, the counts are 68/145 (46.9%) and 37/54 (68.5%), respectively.
+These findings suggest that models may not fully solve the task but instead exploit biased number preferences: notably, o3-Mini—being least likely to guess 7—outperforms the other two o-series models.
+In conclusion, current LLMs struggle with a class of tasks demanding internal object imagination and mental manipulation, relying on an effective working memory.
 6 Conclusion
@@
 We hope these findings encourage future research on integrating working memory mechanisms into LLMs to enhance their reasoning capabilities and human-likeness.
+The Use of Large Language Models
+LLMs were employed in a limited capacity for writing optimization.
+Specifically, the authors provided their own draft text to the LLM, which in turn suggested improvements such as corrections of grammatical errors, clearer phrasing, and removal of non-academic expressions.
+LLMs were also used to inspire possible titles for the paper.
+While the system provided suggestions, the final title was decided and refined by the authors and is not directly taken from any single LLM output.
+In addition, LLMs were used as coding assistants during the implementation phase.
+They provided code completion and debugging suggestions, but all final implementations, experimental design, and validation were carried out and verified by the authors.
+Importantly, LLMs were NOT used for generating research ideas, designing experiments, or searching and reviewing related work.
+All conceptual contributions and experimental designs were fully conceived and executed by the authors.

### v3

@@
 Abstract
-While Large Language Models (LLMs) exhibit remarkable reasoning abilities, we demonstrate that they fundamentally lack a core aspect of human cognition: working memory.
-Human working memory is an active cognitive system that enables not only the temporary storage of information but also its processing and utilization.
-Without working memory, individuals may produce unreal conversations, exhibit self-contradiction, and struggle with tasks requiring mental reasoning.
-Existing evaluations using N-back or context-dependent tasks fail as they allow LLMs to exploit accessible context rather than retain latent information.
-We introduce three novel tasks—(1) Number Guessing, (2) Yes-No Deduction, and Math Magic—that isolate internal representation from external context.
-Across seventeen frontier models spanning four major model families, we consistently observe irrational or contradictory behaviors, highlighting LLMs’ inability to retain and manipulate latent information.
-Our work establishes a new benchmark for evaluating working memory in LLMs and identifies this deficit as a critical obstacle to artificial general intelligence.
+While Large Language Models (LLMs) exhibit remarkable reasoning abilities, we demonstrate that they lack a fundamental aspect of human cognition: working memory.
+Human working memory is an active cognitive system that enables not only the temporary storage of information but also its processing and utilization, enabling coherent reasoning and decision-making.
+Without working memory, individuals may produce unrealistic responses, exhibit self-contradictions, and struggle with tasks that require mental reasoning.
+Existing evaluations using N-back or context-dependent tasks fall short as they allow LLMs to exploit external context rather than retaining the reasoning process in the latent space.
+We introduce three novel tasks: (1) Number Guessing, (2) Yes-No Deduction, and (3) Math Magic, designed to isolate internal representation from external context.
+Across seventeen frontier models spanning four major model families, we consistently observe irrational or contradictory behaviors, indicating LLMs’ inability to retain and manipulate latent information.
+Our work establishes a new benchmark for evaluating working memory in LLMs and highlights this limitation as a key bottleneck for advancing reliable reasoning systems.
 Code and prompts for the experiments are available at https://github.com/penguinnnnn/LLM-Working-Memory.
@@
 1 Introduction
-Large Language Models (LLMs) have demonstrated strong performance across diverse domains, including legal advice Guha et al.
+A notable feature of modern Large Language Models (LLMs) is their ability to perform tasks across a wide range of domains, including law Guha et al.
 (2023), education Wen et al.
@@
 (2024b).
-Beyond task performance, recent studies have examined their intrinsic behaviors, often drawing parallels with human cognition, such as personality Huang et al.
+Most evaluations focus on their extrinsic behaviors—what tasks they can and cannot perform.
+To better understand fundamental, underlying model capabilities, a growing body of work examines their intrinsic behaviors, core abilities that shape the downstream task performance.
+Inspired by cognitive science, recent studies examine whether LLMs exhibit human-like features such as personality Huang et al.
 (2024b), emotion Huang et al.
@@
 (2024b).
-One such behavior is memory, which has attracted increasing attention from both industry and research communities.
-OpenAI was the first to introduce a memory module in ChatGPT (February 2024).
-[^1] https://openai.com/index/memory-and-new-controls-for-chatgpt/ By mid-2025, xAI, Anthropic, and Google integrated memory into Grok, [^2] https://x.com/grok/status/1912670182012801156 Claude, [^3] https://www.anthropic.com/news/claude-4 and Gemini, [^4] https://blog.google/products/gemini/temporary-chats-privacy-controls/ respectively.
-Sam Altman even emphasized that “people want memory,” identifying it as a central feature of the forthcoming GPT-6.555https://futurism.com/disastrous-gpt-5-sam-altman-hyping-up-gpt-6 On the research side, memory modules are also actively studied Zhang et al.
-(2025); Hatalis et al.
-(2023); Xu et al.
-(2025).
+One such ability is memory, which has attracted increasing attention from both industry and research communities.
+OpenAI was the first to introduce a memory module in ChatGPT (February 2024) [^1] https://openai.com/index/memory-and-new-controls-for-chatgpt/ that allowed the model to remember information from previous interactions with a user, such as the user’s facts and preferences.
+The model could access and retrieve these “memories” to be used in later conversations.
+By mid-2025, xAI, Anthropic, and Google have integrated memory into Grok, [^2] https://x.com/grok/status/1912670182012801156 Claude, [^3] https://www.anthropic.com/news/claude-4 and Gemini, [^4] https://blog.google/products/gemini/temporary-chats-privacy-controls/ respectively.
 What is memory?
@@
 These distinctions have also been adopted in machine learning, corresponding to representational learning for raw inputs (sensory memory), in-context computation at test time (working memory), and access to external databases (long-term memory) Weng (2023).
-As a central aspect of human cognition, memory has been widely benchmarked Du et al.
+Researchers in the AI community have investigated long-term memory mechanisms for both individual LLMs Wu et al.
+(2025); Du et al.
+(2025) and LLM agents Zhang et al.
+(2025); Xu et al.
 (2025).
-However, working memory remains relatively underexplored.
+Recent work has shifted from context-dependent approaches (e.g., Chain-of-Thought (CoT) Wei et al.
+(2022) and scratchpads Lanchantin et al.
+(2023)) to explicit storage methods (e.g., text-based Park et al.
+(2023) or vector-based Hatalis et al.
+(2023)) to support lifelong memory Zheng et al.
+(2025); Wang et al.
+(2025b).
+Most studies frame memory as an engineering problem: enabling LLMs to store and retrieve information for later use.
+By contrast, working memory remains relatively underexplored.
 Existing studies often adopt the N-back task Kirchner (1958) to assess LLMs’ working memory Gong et al.
@@
 (2024).
-Yet, a fundamental limitation arises: the critical information for correct responses remains accessible in the model’s input context.
-Unlike humans, who cannot revisit prior steps, LLMs can simply attend to earlier tokens within their context window.
+Yet, a fundamental limitation arises: the critical information for correct responses remains accessible in the model’s input context, allowing models to “look back” rather than actively maintain internal state.
+Therefore, these tests are exploring aspects of the context window, not working memory directly.
+Unlike humans, who cannot revisit prior steps explicitly, LLMs can simply attend to earlier tokens within their context window.
 Fig.
 2 illustrates this discrepancy.
-To more faithfully evaluate working memory, it is necessary to design experiments where the key information is not explicitly present in the context.
+To more faithfully evaluate working memory, it is necessary to design experiments where the key information is not explicitly present in the context and is only available if stored in the working memory of the model.
 Imagine the following scenario: You select a number between one and ten.
 When ready, you are asked, “Is the number greater than five?” If you answer, observers can reasonably infer that the number has entered your conscious awareness (i.e., your working memory), since clear perception is necessary to perform the comparison and provide a response.
-[^6] A possibility remains that the response was given by chance.
-We ask: Do LLMs possess human-like working memory beyond evaluations that merely test their ability to attend to visible information?
-AI models are often viewed as reasoning in two modes: (1) the token space over sequences Wei et al.
+[^5] A possibility remains that your response was given by chance if you tell a lie (Fig.
+1 Case 2) or do not think of a number at all (Fig.
+1 Case 3).
+We ask: Do LLMs possess human-like working memory, or do they only appear to reason by exploiting their context window?
+LLMs are often viewed as reasoning in two modes: (1) the token space over sequences Wei et al.
 (2022); Yao et al.
@@
 (2025).
-Our focus is on the second mode, where evaluating working memory provides insight into whether models can hold and manipulate latent concepts without explicit externalization.
-Success in this capacity could enhance reasoning without reliance on Chain-of-Thought (CoT), as it directly tests the model’s ability to maintain objects and concepts internally.
+We argue that working memory is necessary to enable stronger latent space reasoning, as the model does not have access to its external reasoning tokens.
+Evaluating working memory provides insight into whether models can hold and manipulate latent concepts without explicit externalization.
+Success in this capacity could enhance reasoning without reliance on CoT, as it directly tests the model’s ability to maintain objects and concepts internally.
 Conversely, deficits in working memory impair information processing in humans Gruszka & Nkecka (2017); Cowan (2014), and in LLMs manifest as unrealistic outputs, self-contradictions, and failures on tasks requiring mental manipulation.
-The central challenge in designing such evaluations is epistemic: how can we demonstrate the absence of internal memory when we cannot directly observe a model’s mind?
+The central challenge in designing such evaluations is: How can we demonstrate the presence or absence of internal memory when we cannot directly observe a model’s mind?
 To address this limitation, we design three experiments—(1) Number Guessing, (2) Yes–No Deduction, and (3) Math Magic—that test whether LLMs can internally maintain information that is not explicitly present in the input context.
-Our experiments span 17 state-of-the-art LLMs, both proprietary and open-source, and reveal a consistent inability to perform such tasks.
-The failure is uniform across model families (GPT Hurst et al.
-(2024), Qwen Yang et al.
-(2024a), DeepSeek Liu et al.
-(2024a); Guo et al.
-(2025), LLaMA Grattafiori et al.
-(2024)) and reasoning approaches (CoT Wei et al.
-(2022) or o1-like reasoning Jaech et al.
-(2024)).
-2 Background
+Our experiments span 17 frontier LLMs, both proprietary and open-source, including multiple model families and reasoning approaches.
+Across all settings, the results converge: current LLMs show little evidence of intrinsic working memory.
+Instead, their reasoning appears to depend on externalized context.
+These findings suggest that progress in reasoning will require not only larger models or better prompting, but also closer attention to the mechanisms that could endow LLMs with genuine working memory.
+2 Preliminaries
 Figure 2: a.
@@
 These tasks are widely used because they probe the ability to maintain and manipulate information that is no longer externally visible, thereby capturing the essence of working memory function.
-In the context of LLMs, working memory has been widely discussed.
-However, most studies use the term to describe an LLM’s capacity to process information within a fixed context window Li et al.
+In the context of LLMs, working memory has been used more loosely: most studies use the term to describe an LLM’s capacity to process information within a fixed context window Li et al.
 (2023); Guo et al.
@@
 (2024) evaluate LLMs using N-back tasks.
-As noted in §1, such human-designed tests are not directly valid for LLMs, since models can simply attend to retained tokens in their context window rather than actively maintaining information.
+However, as noted in §1, such human-designed tests are not directly valid for LLMs, since models can simply attend to retained tokens in their context window without actively maintaining information internally.
 2.2 Existing Engineering Solutions
-A separate line of work introduces explicit engineering solutions by equipping LLMs with external memory modules Hu et al.
+A parallel line of work introduces explicit engineering solutions by equipping LLMs with external memory modules Hu et al.
 (2025); Zeng et al.
@@
 Other approaches implement scratch spaces Lanchantin et al.
-(2023) or internal CoT mechanisms Jaech et al.
+(2023), internal CoT mechanisms Jaech et al.
+(2024); OpenAI (2025), and external vector databases Hatalis et al.
 (2023), which may partially mitigate working memory limitations identified in this paper.
 However, these methods do not address whether LLMs possess an intrinsic working memory capacity, analogous to humans’ ability to mentally simulate and manipulate objects.
-The distinction can be illustrated with an analogy: modern calculators make arithmetic trivial, yet schools continue to assess addition, subtraction, multiplication, and division.
+The distinction can be illustrated with an analogy: calculators make arithmetic trivial, yet schools continue to assess addition, subtraction, multiplication, and division.
 The purpose is not solely to solve problems efficiently, but to reveal individuals’ underlying cognitive abilities.
-Similarly, while engineering techniques can extend an LLM’s effective memory, we are ultimately interested in whether the model itself has the basic cognitive capability of working memory.
+Similarly, while engineering techniques can extend an LLM’s effective memory, we are ultimately interested in whether the model intrinsically has the basic cognitive capability of working memory.
 Without it, a model may function adequately through external tools, but its intrinsic reasoning ability remains fundamentally limited.
+2.3 Experimental Design
+All three experiments in this paper are designed to demonstrate that current LLMs lack effective working memory, as evidenced by their inability to perform our proposed tasks.
+Since we cannot let models reveal what they are privately thinking, the three experiments each test a different hypothesis, corresponding to distinct consequences of impaired working memory.
+(1) The Number Guessing game (§3) evaluates whether an LLM’s response distribution across repeated identical queries remains valid.
+(2) The Yes–No Deduction game (§4) examines whether LLMs contradict themselves.
+(3) The Math Magic (§5) assesses whether LLMs’ internal reasoning produces correct outcomes.
+We evaluate 17 frontier LLMs, including GPT (4o Hurst et al.
+(2024) and 4o-Mini OpenAI (2024)), o-series (o1-Mini Jaech et al.
+(2024), o3-Mini, and o4-Mini OpenAI (2025)), LLaMA (3.1 Meta (2024) and 3.3), Qwen-2.5 Yang et al.
+(2024a) (7B and 72B), QwQ Team (2024), and DeepSeek (V3 Liu et al.
+(2024a) and R1 Guo et al.
+(2025)).
+All models are configured with a temperature of $1.0$ and a top‑p value of $1.0$.
 3 Number Guessing
-(a) Probabilities for OpenAI GPT models.
-(b) Probabilities for OpenAI o-series models.
-(c) Probabilities for Meta LLaMA models.
-(d) Probabilities for Qwen and DeepSeek models.
-Figure 3: Probabilities of model answering “Yes” for each number from one to ten.
-Consider a number-guessing game in which a human participant privately selects a number between one and ten, and we ask whether the chosen number is one.
-By repeating this process hundreds of times, we can estimate the probability of the participant selecting the number one.
-Repeating the procedure for each possible number yields an empirical distribution over the participant’s choices.
-It is worth noting that, if the participant actually selects a number and responds honestly, the estimated probabilities across all numbers should sum to one.
-Figure 4: Probabilities of GPT-4o-2024-08-06 answering “Yes” for each number in different ranges.
+Hypothesis.
+Consider a number-guessing game in which a human participant privately selects an integer between one and ten.
+The experimenter then asks whether the chosen number is one.
+By repeating this procedure multiple times, we can estimate the probability of selecting one, denoted $p_{1}$.
+Extending this process to other numbers yields an estimated distribution over all choices, denoted $p_{1},\dots,p_{n}$.
+It is worth noting that, if the participant truly selects a number and responds honestly, the estimated probabilities should form a valid distribution, satisfying $\sum_{i=1}^{n}p_{i}=1$.
+In contrast, if an LLM does not base its responses on an actual hidden choice, the resulting estimates will typically violate this constraint, producing $\sum_{i=1}^{n}p_{i}\neq 1$.
 Setup.
-Leveraging this property, we design an experiment to probe LLMs.
+Leveraging this hypothesis, we design a controlled experiment.
 In each trial, the model is given a fixed prompt: “User: Think of an integer between 1 and 10, but don’t say it to me.
@@
 I’ve thought of an integer between 1 and 10.
-What’s next?” The model is then prompted 200 times for each integer from 1 to 10 with queries such as “Is the number you’re thinking of 7?
-Answer Yes or No.” We record the frequency of “Yes” responses for each number and treat these frequencies as probabilities.
-If the sum of these probabilities deviates significantly from one, it suggests that LLMs either do not think of a specific number or lie to users.
-All models are configured with a temperature of 1.0 and a top‑p value of 1.0.
+What’s next?” The model is then independently prompted 200 times for $i=1,\dots,10$ with queries such as “Is the number you’re thinking of $i$?
+Answer Yes or No.” We record the frequency of “Yes” responses for each number and compute the estimated probabilities $p_{i}$.
+If the sum of these probabilities deviates significantly from one, it suggests that the LLM either is not maintaining a number commitment or lies to users.
 Table 1: The sum of probabilities of each model responding “Yes” for all numbers from one to ten.
@@
 DeepSeek-R1 0.640
-Findings.
+Results.
 Fig.
@@
 Two key observations emerge: (1) Most LLMs never produce a “Yes” response; “No” dominates across models.
-This suggests that LLMs generate unrealistic distributions, implying an inability to produce realistic responses.
+This produces invalid distributions, further indicating that models are estimating the probability of a human guess being correct (10% in our setting, typically very low) rather than maintaining a private number choice.
 Given that LLMs generally follow instructions and do not deliberately deceive, we attribute this behavior to their failure to internally “think of” a number.
-(2) When LLMs do respond affirmatively, they exhibit a marked preference for the number seven.
-This tendency mirrors patterns observed in humans Miller (1956); Kubovy & Psotka (1976).
+(2) When LLMs do respond affirmatively, they exhibit a pronounced preference for the number seven.
+This tendency mirrors human biases Miller (1956); Kubovy & Psotka (1976).
+(a) Probabilities for OpenAI GPT models.
+(b) Probabilities for OpenAI o-series models.
+(c) Probabilities for Meta LLaMA models.
+(d) Probabilities for Qwen and DeepSeek models.
+Figure 3: Probabilities of model answering “Yes” for each number from one to ten.
 We quantify LLM performance on this task using the sum of probabilities.
-A value closer to one indicates better model performs.
+A value closer to one indicates better model performance.
 Table 1 reports these sums for each model.
-Several observations emerge: (1) Newer models do not necessarily outperform older ones.
-Within the GPT family, the 0806 version of GPT-4o (the current target of the “gpt-4o” API) achieves the best performance, surpassing both the 1120 version and GPT-4.1.
+Several observations stand out: (1) Newer models do not necessarily outperform older ones.
+Within the GPT family, the 0806 version of GPT-4o (the model currently served under the “gpt-4o” API) achieves the best performance, surpassing both the 1120 version and GPT-4.1.
 Similarly, LLaMA-3.3 underperforms relative to LLaMA-3.1.
@@
 (3) Overall, LLaMA-3.1 performs best, with the 8B variant outperforming both the 405B and 70B versions.
+Taken together, these findings suggest that acquisition of this capability appears largely stochastic and is less predictable with respect to model scale.
+More broadly, they suggest that the observed memory limitations arise not from model size or training sophistication but from a fundamental architectural deficiency.
+Figure 4: Probabilities of GPT-4o-2024-08-06 answering “Yes” for each number in different ranges.
 Table 2: The sum of probabilities of GPT-4o-2024-08-06 responding “Yes” for all numbers in different ranges.
@@
 40 0.005
-Other Number Ranges.
-We extend our experiments to include a broader range of numbers.
+We further extend our experiments to include a broader range of numbers.
 Given that GPT-4o-2024-08-06 performs best among the OpenAI models, we focus on its behavior across different numerical ranges.
@@
 4.
-In conclusion, LLMs fail to generate correct distributions, suggesting they struggle to produce realistic responses.
-This limitation likely arises from their inability to internally think of numerical values without explicit contextual cues.
+In conclusion, LLMs fail to generate distributions consistent with internally committing to a number.
+Their outputs are either dominated by “No” response or reflect biased heuristics towards seven.
+These findings suggest that LLMs struggle to represent and sustain latent numerical values without explicit contextual grounding, thereby highlighting a gap in working memory-like capacity.
 4 Yes-No Deduction
+Hypothesis.
+“Yes-No” (or the Twenty questions [^6] https://en.wikipedia.org/wiki/Twenty_questions) is a social deduction game commonly used to train human reasoning, classification, and questioning skills.
+In this game, one player privately selects an object, while the opponent asks yes–no questions (e.g., “Is the object heavier than an elephant?”) to progressively narrow down the possibilities and ultimately find the object.
+Consider the decision-making process of the player answering questions, each question requires simply direct comparison between the imagined object and the queried attribute.
+Note that humans typically do not recall all previous questions and answers.
+Instead, they rely on the single reference of the object, without checking self-contradiction with all prior responses.
+We hypothesize that if LLMs cannot maintain such an imagined object in working memory, they can only respond to questions by checking consistency with their prior answers.
+As the number of questions increases, maintaining consistency becomes increasingly difficult, making the task strongly dependent on long-context reasoning.
+To test this, we first instruct the LLM to imagine an object and answer a sequence of comparative questions against the reference objects.
+The goal is to assess whether the model produces self-contradictions.
+For instance, the model might initially answer “Yes” to “Is the object heavier than an elephant?” but later also respond “Yes” to “Is the object lighter than a cat?”, thereby contradicting itself.
 Table 3: Objects ordered by the five properties (smallest to largest).
@@
 Diamond ring
-“Yes-No” is a social deduction game commonly used to train human reasoning, classification, and questioning skills.
-In this game, one player privately selects an object, while the opponent asks yes–no questions (e.g., “Is the object heavier than an elephant?”) to progressively narrow the possibilities and ultimately guess the object.
-Consider the decision-making process of the player answering questions.
-Upon receiving a question, the player simply compares the imagined object with the queried attribute and responds accordingly.
-Note that, humans do not typically recall all previous questions and answers to ensure consistency; instead, they answer each new question based on immediate judgment, without explicitly checking self-contradiction with prior responses.
-We hypothesize that if LLMs lack the working memory to temporarily retain an imagined object, they can only respond to questions by checking consistency with their prior answers.
-As the number of questions increases, maintaining consistency becomes increasingly difficult.
-Consequently, performance on this task is closely tied to the model’s ability for long-context reasoning.
-In this section, we present an experiment in which the LLM is first instructed to imagine an object, followed by a sequence of comparative questions involving that object and other reference objects.
-The goal is to assess whether the model produces self-contradictions when answering sequential questions about the imagined object.
-For instance, the model might initially answer “Yes” to “Is the object heavier than an elephant?” but later also respond “Yes” to “Is the object lighter than a cat?”, thereby contradicting itself.
 Setup.
-We predefined five sets of objects that are commonly regarded as comparable with respect to five properties: volume, length, weight, density, and hardness.
+We predefine five sets of objects that are commonly regarded as comparable with respect to five properties: volume, length, weight, density, and hardness.
 In total, 60 distinct objects are included, as listed in Table 3, ordered by the corresponding property.
 For each question, one property is randomly selected, followed by an object from the corresponding object list.
-The LLM is then prompted to assess whether the object it has in mind is comparative than the given object, where the comparative form can vary in direction (e.g., bigger or smaller for volume).
+The model is then prompted to assess whether the object it imagined is comparative relative to the given object, where the comparative form can vary in direction (e.g., bigger or smaller for volume).
 In each trial, the model is continuously presented with up to 250 such questions.
-We record the number of questions completed before the model exhibits a self-contradiction.
+We record the number of questions it completed before the model exhibits a self-contradiction.
 If no contradiction is observed across all 250 questions, the trial is considered a Pass.
 Each model is tested with 200 trials.
-Table 4: Count of failures of GPT-4o-Mini-2024-07-18 and GPT-4o-2024-08-06.
+Table 4: Count of failures of Yes-No Deduction on the five properties.
 Model Failure V W L D H
@@
 GPT-4o 173 21 42 57 27 26
-Findings.
-Table 4 presents the number of failed trials for the two models.
-The smaller model (GPT-4o-Mini) consistently fails, indicating that it always produces self-contradiction.
-As model capacity increases, GPT-4o successfully passes 27 out of 200 trials.
-This result supports our hypothesis that model performance on this task depends on the their long-context processing ability—an outcome that would be unlikely if the models possessed the working memory for maintaining imagined objects.
-Figure 5 presents histograms of the number of questions each model completes before exhibiting self-contradiction.
-The distribution for GPT-4o-Mini peaks in the 20–30 range, whereas GPT-4o peaks in the 30–40 range.
-Moreover, GPT-4o demonstrates a higher frequency of completions in the 80–130 range compared to GPT-4o-Mini.
-Notably, the types of properties that lead to self-contradictions differ between the two models: GPT-4o-Mini more frequently fails on density and hardness, while GPT-4o shows greater robustness on these attributes.
-Table 5: Ablation studies using GPT-4o-2024-08-06.
+(a) GPT-4o-Mini-2024-07-18 and GPT-4o-2024-08-06.
 Model Failure V W L D H
@@
 Hint + All 145 13 32 34 46 20
-Ablation.
-To ensure that the observed failures are not simply due to LLMs’ inability to rank objects by the five properties, we conduct the following ablation studies: (1) Hint: At the beginning of the prompt, we provide GPT-4o with the object rankings defined in Table 3.
+(b) Ablation studies using GPT-4o-2024-08-06.
+Results.
+Table 4a presents the number of failed trials for the GPT-4o-2024-08-06 Hurst et al.
+(2024) and GPT-4o-Mini-2024-07-18 OpenAI (2024).
+The smaller model (GPT-4o-Mini) consistently fails, while the larger GPT-4o successfully passes 27 out of 200 trials.
+This result supports our hypothesis: model performance on this task depends on their long-context processing ability rather than intrinsic working memory for maintaining imagined objects.
+Figure 5 presents histograms of the number of questions each model completes before exhibiting self-contradiction.
+The distribution for GPT-4o-Mini peaks in the 20–30 range, whereas GPT-4o peaks in the 30–40 range.
+Moreover, GPT-4o demonstrates a higher frequency of completions in the 80–130 range compared to GPT-4o-Mini.
+Notably, the types of properties that lead to self-contradictions differ between the two models: GPT-4o-Mini fails more frequently on density and hardness, while GPT-4o shows greater robustness on these attributes.
+To ensure that the observed failures are not simply due to LLMs’ inability to rank objects by the five properties (i.e., the lack of commonsense knowledge about object properties), we conduct the following ablation studies: (1) Hint: At the beginning of the prompt, we provide GPT-4o with the object rankings defined in Table 3.
 (2) All: For each question, we specify the target object $\mathcal{O}$ by stating explicitly that “the object you are considering is $\mathcal{O}$.” (3) Hint + All: We combine the above two settings.
-Results are shown in Table 5.
-Two key findings emerge: (1) Even when given hints that guarantee knowledge of the correct ranking, GPT-4o still frequently fails, suggesting that the task critically depends on long-context reasoning.
-(2) When the object identity is provided explicitly, error rates decrease, likely because the long-context reasoning task is reduced to a short-context one.
-In conclusion, LLMs exhibit self-contradictions (e.g., claiming an object is both larger than a car and smaller than a soccer ball) as the number of queries increases.
-This behavior suggests their reliance on long-context reasoning rather than a proper working memory.
+Results are shown in Table 4b.
+Two key findings emerge: (1) Providing hints does not prevent contradictions, indicating that the task depends more on long-context reasoning rather than factual knowledge.
+(2) Explicitly specifying the object substantially reduces errors, effectively collapsing the long-context reasoning task into a short-context reasoning problem.
+Across the conditions, models exhibit self-contradictions (e.g., claiming an object is both larger than a car and smaller than a soccer ball) as the number of queries increases.
+This behavior suggests their reliance on long-context reasoning rather than possessing a dedicated working memory for maintaining such an internal state.
 (a) GPT-4o-Mini-2024-07-18.
@@
 5 Math Magic
-Let’s play a game before presenting the final part of our experiments.
+Hypothesis.
+Consider the following recreational arithmetic game, a variant of the Kaprekar routine Kaprekar (1955), which relies on digit manipulation in base 10.
 Think of a three-digit number in which the hundreds and units digits differ (e.g., abc).
@@
 Then, reverse the result—if it is a two-digit number, prepend a zero (e.g., 67 $\rightarrow$ 067 $\rightarrow$ 760).
-Add this reversed number to the previous result What if we can tell that your result is 1089?
-It is math who tells us the secret.
-Such magic, often referred to as “math magic” or “math mentalism,” rely on carefully constructed transformations that yield predictable outcomes, allowing magicians to perform mind-reading.
-To perform this task, humans must internally generate or retain numerical information (e.g., remember a poker card) and manipulate it mentally without external cues.
-This process relies on working memory, a cognitive function that LLMs may lack.
-Without this capacity, LLMs are prone to errors in mathematical tricks, failing to produce the intended outcomes.
-Accordingly, we can design complex math magics to evaluate whether LLMs possess working memory–like capabilities.
+Add this reversed number to the previous result.
+Unsurprisingly, such computation always leads to 1089.
+This deterministic convergence is an example of “mathematical magic” or “math mentalism”—procedures that appear mysterious but are fully explained by the arithmetic structure of decimal digits.
+The invariance arises from the fact that the subtraction step yields a multiple of 99, and the subsequent reversal–addition step collapses all cases to the constant 1089.
+From a cognitive perspective, performing such routines requires humans to encode digits in working memory, apply digit-level transformations (reversing, subtracting, padding), and track intermediate results internally—similar to remembering a poker card during a trick.
+This process relies on working memory, a lack of which would lead to failure to reproduce these deterministic outcomes when asked to simulate the trick.
 Table 5: Operations in the math magic in our experiment. the random number $a$ ranges from 1 to 7, while the random numbers $b$ and $c$ range from 1 to 3.
@@
 Setup.
-Since the 1089 game is well known and likely included in the training data of LLMs, our preliminary experiments show that LLMs can recognize the game and thus invalidate the evaluation.
-Therefore, to more effectively assess LLMs’ multi-step mental manipulation, we select a more complicated math magicbased on the Josephus Problem Schumer (2002).
+Our preliminary experiments show that LLMs can recognize and accurately predict the number 1089, suggesting that this well-known game is likely included in the training data, invalidating this game as an evaluation protocol.
+To more effectively assess LLMs’ capability for multi-step mental manipulation, we select a more complicated routine based on the Josephus Problem Schumer (2002).
 In this task, participants are asked to imagine four numbers and perform a sequence of operations, including duplication, rotation, and removal.
-The procedure is shown in Table 6.
+The full procedure is illustrated in Table 5.
 Ultimately, only two numbers remain, and mathematical constraints guarantee they are identical.
-In our experiment, LLMs are prompted to secretly think of four numbers and mentally execute the sequence of randomized operations.
-We measure the proportion of trials (out of 150) in which the LLM completes the task correctly.
+In our experiment, we prompt LLMs to privately select four numbers and mentally execute the sequence of operations.
+We report the proportion of 150 trials in which the model correctly produced two identical numbers.
+Results.
+Table 6a reports the accuracy of prompting models to output the two numbers directly.
+Most LLMs perform poorly on this task, with notably higher accuracy observed in the LLaMA model family.
+This finding aligns with results from the number guessing game shown in Table 1, where LLaMA models generate more realistic distributions than other models.
+Taken together, these findings point to a consistent trend: while some models perform marginally better, current LLMs generally fail to maintain the internal state required for this kind of sequential manipulation.
 Table 6: LLM performance on the math magic task.
@@
 (b) LLMs with CoT and Large Reasoning Models. GPT-4o-2024-11-20 consistently fails this task.
-Findings.
-Table 7a reports the accuracy of prompting models to output the two numbers directly.
-Most LLMs perform poorly on this task, with notably higher accuracy observed in the LLaMA model family.
-This finding aligns with results from the number guessing game shown in Table 1, where LLaMA models generate more realistic distributions than other models.
-Together, these consistent results suggest that current LLMs generally lack the working memory required for such tasks.
 We further examine whether CoT prompting improves performance on this task.
@@
 Base models prompted to reason step-by-step achieve 10–30% accuracy—substantially higher than without CoT.
-DeepSeek-R1 attains 100% accuracy, and other LRMs also perform well, indicating that extended token-level reasoning substantially improves accuracy.
-However, consistent with our number-guessing experiment, models exhibit a strong preference for the number seven, which appears to influence performance—especially within the o-series.
-Among o1-Mini’s 75 correct predictions, 50 (66.7%) select 7; for o3-Mini and o4-Mini, the counts are 68/145 (46.9%) and 37/54 (68.5%), respectively.
-These findings suggest that models may not fully solve the task but instead exploit biased number preferences: notably, o3-Mini—being least likely to guess 7—outperforms the other two o-series models.
-In conclusion, current LLMs struggle with a class of tasks demanding internal object imagination and mental manipulation, relying on an effective working memory.
-6 Conclusion
-In this study, we present three carefully designed experiments to investigate whether LLMs possess human-like working memory for processing information and generating responses.
+DeepSeek-R1 attains 100% accuracy, and other LRMs also perform well.
+Notably, models also exhibit a strong preference for the number seven, consistent with our number-guessing experiment.
+For example, 66.7% of o1-Mini’s correct predictions, 46.9% of o3-Mini’s, and 68.5% of o4-Mini’s involve the number seven.
+Notably, o3-Mini—being least likely to guess 7—achieves a higher accuracy than other two o-series models.
+These findings suggest that CoT and LRMs can improve accuracy by externalizing intermediate steps, but the success depends on explicit reasoning tokens rather than latent persistence for working memory.
+The persistence of number preference bias and failure on these tasks suggests that current LLMs struggle with tasks that require sustained internal state and mental manipulation.
+6 Discussion
+Summary.
+In this study, we present three carefully designed experiments to investigate whether LLMs exhibit human-like working memory.
 Across all experiments, the results reveal a consistent pattern: LLMs do not exhibit behavior indicative of a functional working memory.
-They fail to internally represent and manipulate transient information across multiple reasoning steps, relying instead on the immediate prompt context.
+They fail to internally represent or manipulate transient information across multiple reasoning steps, relying instead on the immediate prompt context.
 Even advanced prompting strategies, such as CoT prompting, yield only marginal improvements on tasks requiring internal state management.
-This limitation leads to unrealistic responses, self-contradictions, and an inability to perform mental manipulations.
-We hope these findings encourage future research on integrating working memory mechanisms into LLMs to enhance their reasoning capabilities and human-likeness.
-The Use of Large Language Models
-LLMs were employed in a limited capacity for writing optimization.
-Specifically, the authors provided their own draft text to the LLM, which in turn suggested improvements such as corrections of grammatical errors, clearer phrasing, and removal of non-academic expressions.
-LLMs were also used to inspire possible titles for the paper.
-While the system provided suggestions, the final title was decided and refined by the authors and is not directly taken from any single LLM output.
-In addition, LLMs were used as coding assistants during the implementation phase.
-They provided code completion and debugging suggestions, but all final implementations, experimental design, and validation were carried out and verified by the authors.
-Importantly, LLMs were NOT used for generating research ideas, designing experiments, or searching and reviewing related work.
-All conceptual contributions and experimental designs were fully conceived and executed by the authors.
+Implications.
+The absence of working memory manifests in three ways: unrealistic responses, self-contradictions, and inability to perform mental manipulations.
+This deficit directly constrains LLM performance on real-world tasks that require internal state maintenance for execution, including real-world planning tasks such as travel planning Xie et al.
+(2024a); Wang et al.
+(2025a), scientific inquiry Nathani et al.
+(2025), and application navigation Xie et al.
+(2024b); He et al.
+(2024); Lyu et al.
+(2025).
+The challenges are further magnified in multi-agent settings: without working memory, LLM agents quickly lose track in extended dialogues Laban et al.
+(2025), abandon their initial goals (goal drift Arike et al.
+(2025)), or mistakenly adopt others’ perspectives as their own (identity drift Choi et al.
+(2024)).
+Moreover, for LLM-based multi-agent social simulation, the lack of working memory departs LLMs from real-world human subjects, potentially invalidating the simulation as the behavior is fundamentally different Zhou et al.
+(2025).
+In short, the lack of working memory is not just a theoretical concern: it directly undermines reliability, coherence, and validity in applied AI systems.
+In human cognition, both are necessary: we reason aloud and also rely on a silent working memory buffer to hold commitments, track goals, and compare states.
+The absence of this buffer in LLMs may explain why they excel at visible reasoning (e.g., think step by step) yet collapse when asked to “think silently.”
+Future work.
+A natural next step is to explore mechanisms that could grant LLMs intrinsic working memory.
+While engineering approaches such as external text- or vector-based memories can compensate for some deficits, they do not address the core limitation: LLMs’ inability to sustain internal, latent state over time.
+We argue that solutions should move beyond external augmentation toward intrinsic mechanisms—architectural innovations, recurrent depth, or hybrid symbolic–neural components—to provide robust working memory.
+Interpretability studies have shown that specialized attention heads Wang et al.
+(2023); Olsson et al.
+(2022) or expert subnetworks Cai et al.
+(2025) encode distinct functions, hinting at potential internal substrates for working memory.
+Such development could bridge the gap between superficial token recall and genuine state maintenance, enabling more human-like reasoning, advancing both reliability and cognitive plausibility.